OpenStudy (roseliberty):

Please help with calculus, ill medal :)) what is the speed of the ball when it its the ground? D(t) = -t^2+8t+1. d(t) is the height . The maximum height of the ball is17. What is the speed of the ball when it hits the ground, ? time at which ball hits the ground is 8.12 s.

OpenStudy (babykainui):

Hello

OpenStudy (roseliberty):

Hi

OpenStudy (babykainui):

Please notice your equation is of the form \[d(t) = -g t^2+v_0t+h_0\]

OpenStudy (roseliberty):

i tried to substitute it into quadratic formula, which is recommended for this type of problem.

OpenStudy (babykainui):

You can apply the quadratic formula to find the time t

OpenStudy (roseliberty):

i calculated the time is 8.125

OpenStudy (roseliberty):

from the quadratic formula

OpenStudy (jdoe0001):

well, first off, doesn't look like calculus and the ball hits the ground when the y-value is 0, thus, set D(t) to 0 and solve for "x", or get the x-intercepts  \(\bf D(t)=-t^2+8t+1\implies {\color{brown}{ 0}}=-t^2+8t+1\)

OpenStudy (daniel.ohearn1):

Did you try deriving the equation yet?